RC meeting and Run Plan, Thursday March 7
RC Meeting Information
RC Daily Meetings are at 4:30 PM, 2nd floor Counting House Meeting Room
- Bluejeans remote connection information:
To join the Meeting: https://bluejeans.com/861439920 To join via Room System: Video Conferencing System: bjn.vc -or-18.104.22.168 Meeting ID : 861439920 To join via phone : 1) Dial: +1.408.740.7256 (US (San Jose)) +1.888.240.2560 (US Toll Free)) +1.408.317.9253 (US (Primary, San Jose)) (see all numbers - http://bluejeans.com/numbers) 2) Enter Conference ID : 861439920
- Detailed Kinematic Group Settings
- KaonLT-SIDIS Group 2018-2019 Shift Schedule
- Hall-C Controls
- Hall-C Logbook
- MCC White Board
- Energy change in progress. New energy will be 10.219 GeV: https://logbooks.jlab.org/entry/3663391
- Beam expected Friday OWL
- Target operators needed starting Thursday, 2nd half of swing shift. Thursday SWING shift leader on call. All shift workers should report starting Friday OWL.
Current Run Plan
- Beam checkout when beam comes: https://hallcweb.jlab.org/wiki/index.php/Beam_Checkout_Procedures
- Start taking data at setting 33. Keep 3/4 rates in SHMS below 500 kHz, but take maximum available current up to that rate.
Notices in Effect for Shift Crew
- Monitor SHMS 3/4 rates - maximum rate 500 kHz.
Tasks completed during hall access
- Tasks for Wednesday:
- Change HMS polarity (Steve L): DONE
- Increase HMS Q2 current limit to ~ 900 A (Steve L): DONE
- Turn on spectrometer rotation (Steve L): DONE
- Restore HMS Q3 current readback (Steve L): DONE
- HMS Q3 hall probe readback (Ethan): DONE
- Reboot HMS collimator controller (Mark J): DONE
- Reboot HMS Cerenkov readback pressure (Ethan): DONE?
- Change HMS Cerenkov pressure to ~0.45 atm (Howard): DONE
- Update good event counter for CSV (Whit)
- Check trigger and DAQ (Eric P): DONE
- Reboot hccam03: DONE
Planned Accelerator Configuration Changes
- Energy change to 10.219 GeV in progress.